OpenStudy (anonymous):

find the value of x2 +y2 when x=5 and y=3 ??? help anyone explain it too

jimthompson5910 (jim_thompson5910):

x^2+y^2 5^2 + 3^2 25 + 9 34 So x^2+y^2 = 34 when x = 5 and y = 3

hero (hero):

5^2 + 3^2 = 25 + 9 = 34

OpenStudy (anonymous):

ooops typed it wrong guys it was find the value of x2 +y2 when x=-5 and y=3

hero (hero):

same answer

jimthompson5910 (jim_thompson5910):

same answer since (-5)^2 = 25 as well

OpenStudy (anonymous):

really but how about the minus ???

hero (hero):

negative times negative equal positive

OpenStudy (anonymous):

oooooh right :D